OpenStudy (anonymous):

t^2(t+1)-t(2t^2-1) simplify & write in standard form

OpenStudy (mathstudent55):

First, use the distributive property for both sets of parentheses.

OpenStudy (anonymous):

then what? @mathstudent55

OpenStudy (anonymous):

the answer to the question is t^3 +t^2 - t Hope it helped

OpenStudy (anonymous):

I need it explained not just the answer @ajindalaisj

OpenStudy (anonymous):

ok i will elaborate t^2(t+1)-t(2t^2-1) =t^3+t^2-2t^3-t now you combine the like terms =-t^3+t^2-t Hope it helped

OpenStudy (anonymous):

are you sure this is the correct way to work it out? @ajindalaisj

OpenStudy (anonymous):

Yes @pop101 I am 100% sure about that question. Its middle school stuff

OpenStudy (mathstudent55):

\(t^2(t + 1) - t(2t^2 - 1)\) You need to use the distributive property for each set of parentheses. For the first set of parentheses it's pretty straightforward. For the second set of parentheses, remember that the negative sign outside (with the -t) will change all the signs inside the parentheses once you multiply it out. \(t^2(t + 1) - t(2t^2 - 1)\) \(= t^2 * t + t^2 * 1 - t * 2t^2 - t * (-1) \) \(= t^3 + t^2 - 2t^3 + t\) Now we combine like terms: \(= -t^3 + t^2 + t\) It's in standard form because the exponent of t is in descending order.

OpenStudy (mathstudent55):

@ajindalaisj wrote: "the answer to the question is t^3 +t^2 - t Hope it helped ok i will elaborate t^2(t+1)-t(2t^2-1) =t^3+t^2-2t^3-t now you combine the like terms =-t^3+t^2-t Yes pop101 I am 100% sure about that question. Its middle school stuff" The correct answer is: -t^3 + t^2 + t Notice the +t at the end. Be careful with the multiplication of signed numbers. -t * (-1) = t, not -t Also, even a math professor can make a mistake in simple arithmetic, so be careful when you say you're 100% sure and it's simply middle school stuff.